If the wavelength of the fourth harmonic of a standing wave on a string is 7.2 cm, what is the length of the string?
Answer #1

apply V = Lf

f = 330/0.072

f = 4583.33 Hz

apply f = nV/2L

L = 4* 330/4583.33*2)

L = 14.4 cm
Answer #2

Length L = 4*lamdda/2 = 2*7.2 = 14.4 cm